# Personalized Spiking Neural Networks with Ferroelectric Synapses for EEG Signal Processing

### v2

@@ Abstract
Electroencephalography (EEG)-based brain-computer interfaces (BCIs) are strongly affected by non-stationary neural signals that vary across sessions and individuals, limiting the generalization of subject-agnostic models and motivating adaptive and personalized learning on resource-constrained platforms. Programmable memristive hardware offers a promising substrate for such post-deployment adaptation; however, practical realization is challenged by limited weight resolution, device variability, nonlinear programming dynamics, and finite device endurance. In this work, we show that spiking neural networks (SNNs) can be deployed on ferroelectric memristive synaptic devices for adaptive EEG-based motor imagery decoding under realistic device constraints. We fabricate, characterize, and model ferroelectric synapses. We evaluate a convolutional-recurrent SNN architecture under two complementary deployment strategies: (i) device-aware training using a ferroelectric synapse model, and (ii) transfer of software-trained weights followed by low-overhead on-device re-tuning. To enable efficient adaptation, we introduce a device-aware weight-update strategy in which gradient-based updates are accumulated digitally and converted into discrete programming events only when a threshold is exceeded, emulating nonlinear, state-dependent programming dynamics while reducing programming frequency. Both deployment strategies achieve classification performance comparable to state-of-the-art software-based SNNs. Furthermore, subject-specific transfer learning achieved by retraining only the final network layers improves classification accuracy. These results demonstrate that programmable ferroelectric hardware can support robust, low-overhead adaptation in spiking neural networks, opening a practical path toward personalized neuromorphic processing of neural signals.
@@ 1 Introduction
Brain-computer interfaces (BCIs) based on electroencephalography (EEG) have attracted significant interest due to their potential for assistive communication in patients with severe motor impairments [bamdad2015application, kawala2021summary, kanungo2021wheelchair]. From a deployment perspective, practical EEG decoding must operate under stringent constraints on latency and power: decisions should be produced with minimal delay after the user’s intent while maintaining ultra-low energy consumption to ensure acceptable battery lifetime and to limit tissue heating in wearable and implantable settings. More broadly, wearable sensors and implantable devices are increasingly pushing artificial intelligence workloads toward the extreme edge [wang2017flexible, shi2016edge, rancea2024edge], where strict thermal and energy budgets fundamentally shape algorithm and hardware choices [vazquez2021compute]. In this context, EEG-based decoding represents a representative extreme-edge signal-processing workload, requiring both low end-to-end latency and high energy efficiency.
@@
A key challenge in extreme-edge biosignal processing is that conventional frame-based pipelines rely on continuous sampling, digitization, and dense processing, which increase energy consumption in sensing, data transmission, and downstream computation as latency requirements become tighter. In contrast, biological neural systems communicate through sparse, asynchronous events [tayarani2021event], where information is encoded in spike timing and temporal structure, offering an efficient representation for time-series biosignals [garg2021signals]. This event-based sensing and processing paradigm motivates spiking neural networks (SNNs) [maass1997networks], which naturally capture temporal dynamics and provide intrinsic memory through neuronal and synaptic state. Beyond algorithms, brain-inspired computation also motivates architectures that co-localize memory and computation to reduce data movement, an overhead that dominates energy consumption in conventional CMOS processors [horowitz20141]. Neuromorphic engineering seeks to translate these principles into event-driven hardware [mead2020we, burr2017neuromorphic], while in-memory computing aims to reduce data-transfer costs by blurring the boundary between storage and processing [sebastian2020memory].
@@
Within this architectural framework, beyond-CMOS memories have emerged as promising building blocks for in-memory computing [theis2017end, mehonic2024roadmap]. In resistive memory arrays, physical laws such as Ohm’s and Kirchhoff’s laws can be exploited to perform multiply-and-accumulate operations directly within memory, reducing data transfer and its associated energy overhead. Memristive devices can emulate synaptic functionality by encoding network weights as programmable conductance states connecting successive layers. Their nonvolatile retention and sub-femtojoule read energy [covi2022challenges], together with demonstrated multi-level programmability [thomas2024versatile], make them attractive for adaptive processing of non-stationary physiological signals [fang2022compact]. Analog or multi-state behavior suitable for learning has been demonstrated across material systems, including phase-change and valence-change memories [rao2023thousands]. Ferroelectric memristive synapses, which consist of an ultra-thin ferroelectric film (zirconium doped hafnium oxide) between two asymmetric electrodes, further exhibit high endurance (above 110 cycles [begon2022scaled]) and co-integration to CMOS [begon2024back].
@@
However, translating these motivations into practical EEG decoding systems remains challenging. Memristive synapses are constrained by limited effective weight resolution, device-to-device variability, nonlinear and state-dependent programming dynamics, and finite endurance, all of which can degrade accuracy and stability when deploying learned models on hardware. At the system level, two common deployment strategies are (i) offline learning in software followed by weight transfer through device programming, and (ii) direct on-device learning using local update rules [garg2025unsupervised, alibart2013pattern]. Both strategies must contend with the highly non-stationary nature of EEG and the significant variability across trials, recording sessions, and individuals, which limits the generalization of subject-agnostic models and motivates adaptive and personalized learning. At the same time, personalization [zheng2016personalizing] must be achieved under strict constraints on energy consumption, programming frequency, and device lifetime, requiring learning mechanisms that are robust to device non-idealities while enabling low-overhead post-deployment adaptation.
@@
In this work, we investigate the deployment of SNNs for EEG-based motor imagery decoding using ferroelectric memristive synaptic devices. We fabricate, characterize, and model the weight update rule in these devices, with particular emphasis on their nonlinear and state-dependent programming dynamics. Building on this characterization, we introduce a device-aware learning framework in which gradient-based updates are accumulated digitally and converted into discrete programming events only upon crossing a threshold, emulating realistic programming behavior while reducing update activity. Using this framework, we evaluate device-aware on-device learning, demonstrate subject-specific transfer learning through retraining only the final network layers, and study robustness to limited weight resolution and programming variability in weight-transfer scenarios with low-overhead device-aware re-tuning.
@@ 2 Methods
Fig. 1: Spiking neural network architecture. The time series signals from all the 64 electrodes were used where the participants imagine limb movements. The electrode potential at each time step is converted to a 2D map, which is passed to the network from ref. [kumar2022decoding] comprising four convolutional layers followed by a recurrent and two fully connected layers.
@@ 2.1 Motor imagery dataset
In this study, we used the Physionet [goldberger2000physiobank] motor imagery dataset [schalk2004bci2000]. This database comprises recording from 109 participants through a 64 channel EEG headset using BCI2000 system with a sampling rate of 160 Hz. These recordings were obtained from participants engaged in a sequence of motor and imagery activities. The tasks related to action are not considered in this study, because the model should be trained on people not able to perform actions. Same as the study by [kumar2022decoding], data from six subjects were removed, resulting in 103 subjects used in our study. The recordings are classified in two subsets of imagined actions and left hand imagery task contained 2336 trials in total and right hand imagery task contained 2299 trials in total. The signals were band-pass filtered between 0.1-80 Hz, and a 1-second segment (0-1 s after cue) was extracted for each trial.
@@ 2.2 Spiking neural network
Our model follows the spiking neural network (SNN) architecture proposed in [kumar2022decoding]. We preserve the overall topology, including the spike-encoding stage, a stack of three convolutional spiking layers (CONV1–CONV3), and two fully connected layers (FC1–FC2) for classification. The signals from the 64 EEG channels are projected onto a spatial layout ($10\times 11$), after which CONV1 extracts local spatial features. Conv2 and Conv3 extract higher-level spatial features at each time step. A temporal convolution layer (TC1) then combines the activity from the last three time steps, capturing short-range temporal patterns before the recurrent and fully connected layers. At each time step, the network produces a vector of classification scores. These logits are modulated by a set of learnable temporal weights $\mathbf{w}_{ts}$ and subsequently aggregated to yield the final prediction:
@@ 2.4 Synaptic device
A ferroelectric synaptic device with the same functional stack as that of [begon2024back] was fabricated on a silicon subtrate. The substrate is a 0.8 µm thick thermal SiO2 on Si chip. Using Atomic Layer Deposition (ALD), 20 nm of TiN was deposited at 300°C, then 45 cycles of WOx at 360°C. Afterwards, a [HfO2/ZrO2] nanolaminate is deposited at 300°C consisting of five supercycles, each comprising five cycles with tetrakis (ethylmethylamino) hafnium (IV) and O2, and ten cycles with bis (methylcyclopentadienyl) (methyl) (methoxy) zirconium (IV) and O2. Ten nanometers of TiN were deposited in situ. Crystallization was performed with a millisecond flash lamp anneal: the sample was preheated at 450°C, followed by a 20 ms pulse of 90 J/cm2. A 50 nm thick W metal electrode was then sputtered. The top electrode was defined by UV lithography and reactive ion etching (RIE). A 100 nm thick SiO2 passivation layer was then sputtered. Vias to the device’s top electrode contacts were defined by UV lithography. The SiO2 layer was etched by RIE. A cross-section is represented in Fig. 2(a).
@@
Fig. 2: Ferroelectric synaptic device programming. a Schematic of the ferroelectric synaptic device stack and the programming scheme, consisting of write pulses with fixed pulse width (50 µs) and increasing amplitudes ($V_{\mathrm{prog}}$). Positive-polarity pulses induce long-term depression (LTD), followed by negative-polarity pulses that induce long-term potentiation (LTP). The device conductance is read after each programming pulse. b Characterization results showing the applied programming-pulse amplitudes (top) and the corresponding evolution of device conductance over time (bottom) for ten programming cycles. Programmed conductance values grouped by pulse amplitude and fitted with Gaussian distributions for LTD (c) and LTP (d) pulse sequences. e Standard deviation of the fitted conductance distributions plotted as a function of the programmed conductance level.
@@
To characterize multiple conductance states, a 8300 µm2 device was repeatedly cycled between high and low-resistance states (HRS/LRS) applying a sequence of programming pulses, as shown in Fig. 2(b). The programming pulses had a constant width of 50 µs and their amplitude gradually increased from 0 V to +1 V for potentiation and from 0 V to -1.25 V for depression, in +/-50 mV steps. For read operations, bipolar triangular pulses of +/-100 mV were applied, and the resulting current was measured using a source-measure unit (SMU). The measured resistance values ranged from approximately 10 MΩ to 100 MΩ. The physical mechanisms leading to a change in conductance upon programming pulses are detailed in ref.[begon2024back]: at each pulse, the direction of the polarization for a fraction of the ferroelectric domains flips. The screening of the polarization results in a local redistribution of charges, modifying the conduction through the insulator.
@@
The programmable conductance states were grouped according to the magnitude of the programming pulses, as shown in Fig. 2(c-d) for long-term depression (LTD) and long-term potentiation (LTP).
@@ 2.5 Weight mapping
such that the trained weights typically satisfy $w_{\mathrm{SNN}}\in[-\mathrm{bound},\mathrm{bound}]$, consistent with the fan-in-based Kaiming He weight initialization scheme commonly used in deep neural networks [he2015delving]. During all the experiments including the noise addition and on-device learning, the weights were clamped to this range. These weights were then rescaled into a differential weight
@@
Bipolar weights are represented using a differential-pair scheme in which the complementary device is fixed at mid-level conductance, while the active device alone encodes the signed weight. Based on retention measurements reported in [begon2024back], the programmed mid-level conductance state was found to be stable over time. Consequently, the hardware update model is applied only to the active device. This differs from conventional differential encoding, commonly used in phase change memory (PCM) [pistolesi2024differential] and other device technologies, where one device stores positive and the other negative weights, motivated by highly asymmetric potentiation and depression characteristics. In contrast, ferroelectric devices exhibit granular and bidirectional conductance change, enabling both LTP and LTD on a single device. This eliminates conductance saturation and reduces refresh operations typically required in two-device differential schemes [nandakumar2020mixed]. Also, this scheme allows to use only one column of devices as a reference column and not duplicate the full array. Offset-based bipolar synapses have also been demonstrated in hardware using digital offset correction [meng2021digital], in online learning with magnetic tunnel junctions [mondal2019situ], and through reference columns to mitigate sneak-path effects [fouda2019mask].

### v3

@@ Abstract
Electroencephalography (EEG)-based brain-computer interfaces (BCIs) are strongly affected by non-stationary neural signals that vary across sessions and individuals, limiting the generalization of subject-agnostic models and motivating adaptive and personalized learning on resource-constrained platforms. Programmable memristive hardware offers a promising substrate for such post-deployment adaptation; however, practical realization is challenged by limited weight resolution, device variability, nonlinear programming dynamics, and finite device endurance. In this work, we show that spiking neural networks (SNNs) can be deployed on ferroelectric memristive synaptic devices for adaptive EEG-based motor imagery decoding under realistic device constraints, achieving classification performance comparable to software-based SNNs. We fabricate, characterize, and model the weight update in ferroelectric synapses. We then evaluate the deployment of convolutional-recurrent SNN architecture using two strategies. First, we adapt to SNNs a mixed precision strategy in which gradient-based updates are accumulated digitally and converted into discrete programming events only when a threshold is exceeded. Additionally, the weight update is device-aware and accounts for the nonlinear, state-dependent programming dynamics. During learning and adaptation, this scheme mitigates possible endurance and energy constraints. Second, we evaluate the transfer of software-trained weights followed by low-overhead on-device re-tuning. We show that, subject-specific transfer learning achieved by retraining only the final network layers improves classification accuracy. These results demonstrate that programmable ferroelectric hardware can support robust, low-overhead adaptation in spiking neural networks, opening a practical path toward personalized neuromorphic processing of neural signals.
@@ 1 Introduction
Brain-computer interfaces (BCIs) based on electroencephalography (EEG) have attracted significant interest due to their potential for assistive communication in patients with severe motor impairments [4, 31, 30]. From a deployment perspective, practical EEG decoding must operate under stringent constraints on latency and power: decisions should be produced with minimal delay after the user’s intent while maintaining ultra-low energy consumption to ensure acceptable battery lifetime and to limit tissue heating in wearable and implantable settings. More broadly, wearable sensors and implantable devices are increasingly pushing artificial intelligence workloads toward the extreme edge [63, 52, 44], where strict thermal and energy budgets fundamentally shape algorithm and hardware choices [61]. In this context, EEG-based decoding represents a representative extreme-edge signal-processing workload, requiring both low end-to-end latency and high energy efficiency.
@@
A key challenge in extreme-edge biosignal processing is that conventional frame-based pipelines rely on continuous sampling, digitization, and dense processing, which increase energy consumption in sensing, data transmission, and downstream computation as latency requirements become tighter. In contrast, biological neural systems communicate through sparse, asynchronous events [57], where information is encoded in spike timing and temporal structure, offering an efficient representation for time-series biosignals [18]. This event-based sensing and processing paradigm motivates spiking neural networks (SNNs) [36], which naturally capture temporal dynamics and provide intrinsic memory through neuronal and synaptic state. Beyond algorithms, brain-inspired computation also motivates architectures that co-localize memory and computation to reduce data movement, an overhead that dominates energy consumption in conventional CMOS processors [28]. Neuromorphic engineering seeks to translate these principles into event-driven hardware [37, 10], while in-memory computing aims to reduce data-transfer costs by blurring the boundary between storage and processing [50].
@@
Within this architectural framework, beyond-CMOS memories have emerged as promising building blocks for in-memory computing [58, 38]. In resistive memory arrays, physical laws such as Ohm’s and Kirchhoff’s laws can be exploited to perform multiply-and-accumulate operations directly within memory, reducing data transfer and its associated energy overhead. Memristive devices can emulate synaptic functionality by encoding network weights as programmable conductance states connecting successive layers. Their nonvolatile retention and sub-femtojoule read energy [12], together with demonstrated multi-level programmability [59], make them attractive for adaptive processing of non-stationary physiological signals [15]. Analog or multi-state behavior suitable for learning has been demonstrated across material systems, including phase-change, valence-change memories [45] and CMOS-compatible ferroelectric resistive weights [8]. Ferroelectric memristive synapses consist of an ultra-thin ferroelectric film (zirconium doped hafnium oxide) between two asymmetric electrodes. They are ideal candidates for neuromorphic hardware for four main reasons: programming is ultrafast, with low energy (allegedly below one nanosecond [3] and one pJ [7]) and high endurance (above 110 cycles [6]). In contrast, memristive technologies relying on ion motion or a change of phase typically require 10 pJ and can only endure 106 cycles [14]. Ferroelectric memristive weights can be read non-destructively at low voltages ($100~\mathrm{mV}$), with fJ energy [5]. Finally, the ferroelectric switching dynamics mimic the biological, synaptic plasticity [9]. Ferroelectric materials find applications beyond the field of neuromorphic hardware [40], for example as non-volatile memories [53], photodetectors [26], energy harvesting devices [47].
@@
However, translating these motivations into practical EEG decoding systems remains challenging. Memristive synapses are constrained by limited effective weight resolution, device-to-device variability, nonlinear and state-dependent programming dynamics, and finite endurance, all of which can degrade accuracy and stability when deploying learned models on hardware. At the system level, two common deployment strategies are (i) offline learning in software followed by weight transfer through device programming, and (ii) direct on-device learning using local update rules [19, 2]. Both strategies must contend with the highly non-stationary nature of EEG and the significant variability across trials, recording sessions, and individuals, which limits the generalization of subject-agnostic models and motivates adaptive and personalized learning. At the same time, personalization [65] must be achieved under strict constraints on energy consumption, programming frequency, and device lifetime, requiring learning mechanisms that are robust to device non-idealities while enabling low-overhead post-deployment adaptation.
@@
In this work, we investigate the deployment of SNNs for EEG-based motor imagery decoding using ferroelectric memristive synaptic devices. The chosen technology is CMOS compatible: in ref. [8], NMOS selectors were used to partially or totally prevent programming in ”1T-1R” cells (1 transistor - 1 resistive weight). We fabricate, characterize, and model the weight update rule in these devices, with particular emphasis on their nonlinear and state-dependent programming dynamics. Building on this characterization, we introduce a device-aware learning framework in which gradient-based updates are accumulated digitally and converted into discrete programming events only upon crossing a threshold, emulating realistic programming behavior while reducing update activity. Using this framework, we evaluate device-aware on-device learning, demonstrate subject-specific transfer learning through retraining only the final network layers, and study robustness to limited weight resolution and programming variability in weight-transfer scenarios with low-overhead device-aware re-tuning.
@@ 2 Methods
Fig. 1: Spiking neural network architecture. The time series signals from all the 64 electrodes were used where the participants imagine limb movements. The electrode potential at each time step is converted to a 2D map, which is passed to the network from ref. [32] comprising four convolutional layers followed by a recurrent and two fully connected layers.
@@ 2.1 Motor imagery dataset
In this study, we used the Physionet [25] motor imagery dataset [49]. This database comprises recording from 109 participants through a 64 channel EEG headset using BCI2000 system with a sampling rate of 160 Hz. These recordings were obtained from participants engaged in a sequence of motor and imagery activities. The tasks related to action are not considered in this study, because the model should be trained on people not able to perform actions. Same as the study by [32], data from six subjects were removed, resulting in 103 subjects used in our study. The recordings are classified in two subsets of imagined actions and left hand imagery task contained 2336 trials in total and right hand imagery task contained 2299 trials in total. The signals were band-pass filtered between 0.1-80 Hz, and a 1-second segment (0-1 s after cue) was extracted for each trial.
@@ 2.2 Spiking neural network
Our model follows the spiking neural network (SNN) architecture proposed in [32], which we adopt because it was specifically developed for EEG motor-imagery decoding and provides a suitable balance between spatial feature extraction, temporal modeling, and network complexity. We preserve the overall topology, including the spike-encoding stage, three convolutional spiking layers (CONV1–CONV3), a temporal convolution layer, a recurrent layer, and two fully connected layers (FC1–FC2) for classification. The 64 EEG channels are projected onto a $10\times 11$ spatial grid with zero padding at unoccupied locations, yielding a 2D representation that preserves the approximate scalp topography and local electrode neighborhood relationships, consistent with prior EEG spatial embedding approaches [11] and standardized electrode placement [1]. The grid accommodates the full electrode arrangement while retaining empty positions where no electrode is present, thereby preserving relative spatial structure rather than forcing the channels into an artificial dense image.
Within this representation, the use of $3\times 3$ convolutional kernels allows the network to capture local spatial correlations among neighboring electrodes without introducing an excessively large number of parameters. Stacking multiple such layers progressively enlarges the effective receptive field, so that CONV1 extracts local spatial patterns, while CONV2 and CONV3 build progressively more abstract spatial features at each time step. All three convolutional layers use $3\times 3$ kernels, which provide a good compromise between capturing short-range spatial correlations and limiting parameter count; stacking such layers also progressively enlarges the effective receptive field, following common deep CNN design principles [54]. In parallel, the number of feature maps increases with depth (64, 128, and 256 for CONV1–CONV3, respectively), forming a hierarchical or pyramidal feature extractor in which earlier layers capture simpler local structure and deeper layers encode richer spatial combinations [60]. After the spatial feature extractor, the temporal convolution layer (TC1) uses a kernel spanning three time steps to capture short-range temporal dynamics, while the recurrent layer models longer-range temporal dependencies across the sequence.
At each time step, the network produces a vector of classification scores. These logits are modulated by a set of learnable temporal weights $\mathbf{w}_{ts}$ and subsequently aggregated to yield the final prediction:
@@
These neurons are modeled using a discrete-time LIF neuron formulation. This choice is motivated by the fact that the EEG signals are temporal data, and LIF neurons intrinsically model the temporal dynamics of the signal. In addition, EEG inputs are acquired as sampled time series and are therefore naturally processed in discrete time steps, whereas surrogate-gradient training requires explicit time unrolling of neuronal dynamics during backpropagation through time.
@@ 2.4 Synaptic device
A ferroelectric synaptic device with the same functional stack as that of [8] was fabricated on a silicon subtrate. The substrate is a 0.8 µm thick thermal SiO2 on Si chip. Using Atomic Layer Deposition (ALD), 20 nm of TiN was deposited at 300°C, then 45 cycles of WOx at 360°C. Afterwards, a [HfO2/ZrO2] nanolaminate is deposited at 300°C consisting of five supercycles, each comprising five cycles with tetrakis (ethylmethylamino) hafnium (IV) and O2, and ten cycles with bis (methylcyclopentadienyl) (methyl) (methoxy) zirconium (IV) and O2. Ten nanometers of TiN were deposited in situ. The ALD deposition was followed by a 50 nm of tungsten sputtering. Crystallization was performed with a millisecond flash lamp anneal: the sample was preheated at 450°C for 120 s, followed by a 20 ms pulse of 90 J/cm2. The top electrode was defined by UV lithography and reactive ion etching (RIE). A 100 nm thick SiO2 passivation layer was then sputtered. A cross-section is represented in Fig. 2(a). -
@@
Fig. 2: Ferroelectric synaptic device programming. a Schematic of the ferroelectric synaptic device stack and the programming scheme, consisting of write pulses with fixed pulse width (50 µs) and increasing amplitudes ($V_{\mathrm{prog}}$). Positive-polarity pulses induce long-term depression (LTD), followed by negative-polarity pulses that induce long-term potentiation (LTP). The device conductance is read after each programming pulse. b Characterization results showing the applied programming-pulse amplitudes (top) and the corresponding evolution of device conductance over time (bottom) for ten programming cycles. c Current density–voltage (J-V) characteristics of the synaptic device, showing the forward and reverse sweep directions. Programmed conductance values grouped by pulse amplitude and fitted with Gaussian distributions for LTD (d) and LTP (e) pulse sequences. f Standard deviation of the fitted conductance distributions plotted as a function of the programmed conductance level.
@@
To characterize multiple conductance states, a 8300 µm2 device was repeatedly cycled between high and low-resistance states (HRS/LRS) applying a sequence of programming pulses, as shown in Fig. 2(b). The programming pulses had a constant width of 500 µs and their amplitude gradually increased from 0 V to +1 V for potentiation and from 0 V to -1.25 V for depression, in +/-50 mV steps. For read operations, bipolar triangular pulses of +/-100 mV were applied, and the resulting current was measured using a source-measure unit (SMU). The measured resistance values ranged from approximately 10 MΩ to 100 MΩ. The physical mechanisms leading to a change in conductance upon programming pulses are detailed in ref.[8]: at each pulse, the direction of the polarization for a fraction of the ferroelectric domains flips. The screening of the polarization results in a local redistribution of charges, modifying the conduction through the insulator. In addition, a current density–voltage (J-V) sweep was performed to obtain the characteristic hysteresis shown in Fig. 2(c).
@@
The programmable conductance states were grouped according to the magnitude of the programming pulses, as shown in Fig. 2(d-e) for long-term depression (LTD) and long-term potentiation (LTP).
@@ 2.5 Weight mapping
such that the trained weights typically satisfy $w_{\mathrm{SNN}}\in[-\mathrm{bound},\mathrm{bound}]$, consistent with the fan-in-based Kaiming He weight initialization scheme commonly used in deep neural networks [27]. During all the experiments including the noise addition and on-device learning, the weights were clamped to this range. These weights were then rescaled into a differential weight
@@
Bipolar weights are represented using a differential-pair scheme in which the complementary device is fixed at mid-level conductance, while the active device alone encodes the signed weight. Based on retention measurements reported in [8], the programmed mid-level conductance state was found to be stable over time. Consequently, the hardware update model is applied only to the active device.